OpenStudy (anonymous):

what is an example of an extraneous solution? the problem im working on is: radical x+3+4=6

OpenStudy (jdoe0001):

\(\bf \sqrt{x}+3+4 = 6\quad ?\)

OpenStudy (anonymous):

yes lol

OpenStudy (jdoe0001):

\(\bf \sqrt{x}+3+4 = 6 \implies \sqrt{x} = -1\) I gather some may say that's extraneous, since a square root wouldn't give a negative value as answer

OpenStudy (anonymous):

you just saved my life lol